OpenStudy (anonymous):

convert to logarithmic equation 5\[^{-3}\]= 1/125

OpenStudy (anonymous):

5^{-3}\] =1/125

OpenStudy (anonymous):

5^{-3}\] =1/125

OpenStudy (anonymous):

its not coming out right

OpenStudy (a_clan):

right

OpenStudy (anonymous):

5^-3=1/125

OpenStudy (anonymous):

I'm not sure how they want you to express it as a logarithmic equation..

OpenStudy (a_clan):

log ( 1/125) = -3 5

OpenStudy (a_clan):

Because when you raise 5 to the power -3 , you get 1/125

OpenStudy (amistre64):

logs are simply an expression of exponents. \(5^{-3}=\frac{1}{125} \iff \left(\frac{1}{5}\right)^3=\frac{1}{125}\)

OpenStudy (amistre64):

\[log_{1/5}(3)=1/125\] is the best I can make out of it, but I cant recall if we are allowed to have a rational base, cant see any reason whynot tho

OpenStudy (amistre64):

got that backwards...

OpenStudy (amistre64):

\[log_{1/5}(1/125)=3\]

OpenStudy (amistre64):

or just as well; \[log_{5}(1/125)=-3\]

OpenStudy (anonymous):

Sure, or \[-3log(5) = log(1) - log(125)\] Or all kinds of other options..